OpenStudy (anonymous):

What is the 35th term of the arithmetic sequence where a1 = 13 and a17 = –67?

OpenStudy (shubhamsrg):

you want your AP home work to be done online isnt it ?

OpenStudy (anonymous):

its not AP?

OpenStudy (shubhamsrg):

yes its AP,,ofcorse it is,,mentioned in the ques..

OpenStudy (shubhamsrg):

well using a1 and a17,,wont you be able to calculate common difference? use this : a_n = a + (n-1)d

OpenStudy (anonymous):

\[a_{1}=13\] \[a_{17}=a+16d=-67\] \[13+16d=-67\] \[16d=-80\] \[d=-5\] hence \[a_{35}=13+34(-5)=-157\]